OpenStudy (confusionist):

Can someone help me solve this? cos(x^2+45)+tan(4x+90) = sin(3x+120) A step by step would be great.

OpenStudy (anonymous):

by solving do you mean proving?

OpenStudy (confusionist):

Yes. :)

myininaya (myininaya):

What methods have been discussed?

myininaya (myininaya):

We are solving for x not proving this. This isn't an identity.

myininaya (myininaya):

Or does it say to disprove or prove?

OpenStudy (confusionist):

Ah yes, my bad. It says solve.

myininaya (myininaya):

Can you tell me what methods in class have been discussed?

OpenStudy (confusionist):

None, this is from a book with only problems I picked up at barnes and noble. We haven't done this and wont do it for another year or two. I've had no problems with the book up to this point.

myininaya (myininaya):

can you tell me what methods the book has discussed?

OpenStudy (confusionist):

None. It's a book of only problems. It starts from Geometry and works it's way up in problems.

myininaya (myininaya):

Name of book please

OpenStudy (confusionist):

Math Review for the High School Classroom

OpenStudy (confusionist):

I looked for it online and can't find it.

myininaya (myininaya):

I didn't see the book. :( So there are no examples?

OpenStudy (confusionist):

No examples, just questions. :c How did you learn it?

myininaya (myininaya):

Honestly, I don't know without seeing some example or them mentioning some method they want to use.

OpenStudy (confusionist):

Ah, okay. Thanks, anyway. :)

OpenStudy (shamil98):

cos(x^2+45)+tan(4x+90) = sin(3x+120) cos(x^2+45)+sin/cos(4x+90) = sin(3x+120) dividing everything by sin. tan(x^2+45) + sec(4x+90) = 3x + 120 i don't remember trig that well, but you can convert it to this i think.

myininaya (myininaya):

You can't do that... Like you can't divide by the trig part The trig part is a function not a number or variable on its own

OpenStudy (shamil98):

oh

OpenStudy (primeralph):

Are the angles in radians or degrees?

OpenStudy (primeralph):

I'll assume degrees for now.

OpenStudy (confusionist):

Degrees, yeah. ^^

hartnn (hartnn):

45,90,120 suggest they are in degrees... we can convert sin and cos in tan, like cos A = tan (sqrt (1-A^2)/A) something like this and then use tan A+tan B formula...... but if thats possible, the algebra is going to be very ugly....

myininaya (myininaya):

but @hartnn tan(90) does not exist you were talking about the sum identity for tan right?

OpenStudy (primeralph):

@myininaya tan(a+90) = -1/tan(a), since they're perpendicular.

hartnn (hartnn):

not tan(A+B) tan A +tan B , where B = 4x+90 and yeah, tan (A+90) = - cot 90 so maybe convert cos into cot and use cotA-cotB formula... i don't think that will be of any use though....

OpenStudy (primeralph):